OpenStudy (anonymous):

Write the equation x – 3y = 6 in slope intercept form. Show your work.

OpenStudy (mathstudent55):

Slope intercept form is: y = mx + b You need to solve your equation for y.

OpenStudy (mathstudent55):

Start by subtracting x from both sides. Then divide both sides by -3.

OpenStudy (anonymous):

idk how to find x

OpenStudy (mathstudent55):

You don't need to find x. Just start by subtracting x from both sides: x - 3y = 6 -x -x ----------------- What do you get?

OpenStudy (anonymous):

6x

OpenStudy (mathstudent55):

You can't combine a term with x and a term without x. On the left side, you have x - 3y Whan you subtract x from the left you have: x - x - 3y You can combine the two terms with x together: x - x = 0, so you end up wityh only -3y on the left side. On the right side, you have 6. Now you subtract x, so you end up with 6 - x which is the same as -x + 6. x - 3y = 6 -x -x ----------------- -3y = -x + 6 Now you need to divide both sides by -3 to end up with y by itself.

OpenStudy (anonymous):

1=-2 ?????????

OpenStudy (anonymous):

?

OpenStudy (mathstudent55):

Now you divide both sides by -3 to get rid of the -3 multiplying the y. -3y/(-3) = -x/(-3) + 6/(-3) \(y = \frac{1}{3}x - 2\)

OpenStudy (anonymous):

sooooooooooo y=1/2x-2 is the slope intercept form of my original question lol???

OpenStudy (mathstudent55):

Yes, that is it. It is in y = mx + b form. In this case, m = 1/2 and b = -2

OpenStudy (anonymous):

thank youuuuuuuuuuuuuuuuuu!!!!!!!!!!!!!!!!!!!

OpenStudy (mathstudent55):

weeeeeeeeeeeellllllllllllcccccccccccccccccccmmmmmmmmmmmmmmm!

OpenStudy (anonymous):

\(^~^)/